# IFJPAN-IV-2023-3

### v1

@@ 2.1 The $f_{i}\bar{f}_{i}\to\tau^{-}\tau^{+}$, $i=e,u,d$ case
$R_{33}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{2}}{8N_{i}D_{Z}(s)}\big{\{}16\beta\gamma^{2}\text{cos}(\theta)a_{\tau}\,a_{i}\,v_{\tau}\,v_{i}+\left(\gamma^{2}-1\right)$
@@
$R_{14}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{3}}{2N_{i}D_{Z}(s)}\text{sin}(\theta)\big{\{}\text{Im}(Y(s))\bigl{[}\beta\,\gamma^{2}\,v_{\tau}\,\text{cos}(\theta)(a_{i}^{2}+v_{i}^{2})$
@@
$R_{24}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{3}}{2N_{i}D_{Z}(s)}\text{sin}(\theta)\big{\{}\text{Im}(X(s))\bigl{[}\left(\gamma^{2}-1\right)v_{\tau}\,\text{cos}(\theta)(a_{i}^{2}+v_{i}^{2})$
@@
$R_{34}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{2}}{4N_{i}D_{Z}(s)}\big{\{}-4a_{i}\,v_{i}\,\text{cos}(\theta)(\left(\gamma^{2}-1\right)a_{\tau}^{2}+\gamma^{2}v_{\tau}^{2})$
@@
$R_{44}^{(Z)}=\frac{g_{Z}^{4}m^{4}\gamma^{2}}{8N_{i}D_{Z}(s)}\big{\{}16\beta\gamma^{2}\text{cos}(\theta)\,a_{\tau}\,a_{i}\,v_{\tau}\,v_{i}+\left(\gamma^{2}-1\right)(3+\text{cos}(2\theta))$
@@
  • Finally, the longitudinal polarization ${\cal P}_{L}=r_{34/43}$ of the $\tau^{-}$ and $\tau^{+}$ in the SM arises due to the vector coupling $v_{\tau}$. This is well known and ${\cal P}_{L}$ was measured already at LEP 1, as noted in Ref. [28]. The longitudinal polarization, as can be seen from Eqs. (21), depends on the weak anomalous magnetic moment ${\rm Re}(X)$ as well.
@@
We would like to acknowledge that in case of $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process, the terms linear in the spins and dipole moments were already derived in Refs. [27, 22]. However, our results represent some important extensions: they are obtained for various initial state fermions $i=(e,\,u,\,d)$, include the complete spin-correlation matrix between the decaying $\tau^{-}$ and $\tau^{+}$ leptons, and also include $Z$-boson exchange, $\gamma^{*}$ exchange and $\gamma^{*}\,Z$ interference.
@@ 2.2 The $f_{i}f_{i}\to\tau^{-}\tau^{+}$, $i=\gamma$ case
$R^{{\gamma\gamma}}_{13}$ $=$ $R^{\gamma\gamma}_{31}=\frac{e^{4}}{2D^{2}}\,\gamma\,\beta^{2}\big{[}(\beta^{2}+A\,(\beta^{2}-2)-1)\cos(2\theta)+A\beta^{2}-\beta^{2}+1\big{]}\,|\sin(2\theta)|,$
@@
$R^{{\gamma\gamma}}_{23}$ $=$ $-R^{\gamma\gamma}_{32}=\frac{e^{4}B}{2D^{2}}\,\gamma\,\beta\,(\beta^{2}\cos(2\theta)-3\beta^{2}+2)\,|\sin(2\theta)|,$
@@
These elements satisfy the condition $R^{{\gamma\gamma}}_{i,j}(\theta)=R^{{\gamma\gamma}}_{i,j}(\pi-\theta)$ for $0\leq\theta\leq\pi$, as follows from identity of the photons.
@@
$+2(s^{-}_{2}s^{+}_{3}-s^{-}_{3}s^{+}_{2})\,\gamma(\beta^{2}\cos(2\theta)-3\beta^{2}+2)\,|\sin(2\theta)|\Big{]}.$
@@ 3 Transverse spin correlations of elementary processes
In this Section, we present numerical results for $R_{i,j}$ for the process $f_{i}\,\bar{f}_{i}\to\tau^{-}\tau^{+}$ obtained with analytical formulas of Section 2, as a function of the invariant mass of $\tau$-lepton pair.
@@
Figure 1: Transverse spin-correlation elements of the $r_{i,j}$ matrix: $r_{11}$ (solid lines) and $r_{22}$ (dashed lines) for the electron-positron initial state. The angle $\theta$ is chosen $\pi/3$ in the left plot and $32\pi/$ 3 in the right plot. The effective couplings of $Z$ to leptons from Table 2 are used. Dipole moments are not included.
Figure 1: Transverse spin-correlation elements of the $r_{i,j}$ matrix: $r_{11}$ (solid lines) and $r_{22}$ (dashed lines) for the electron-positron initial state. The angle $\theta$ is chosen $\pi/3$ in the left plot and $32\pi/$ 3 in the right plot. The effective couplings of $Z$ to leptons from Table 2 are used. Dipole moments are not included.
@@
We have also used distributions from Fig. 2 to test our formulas with respect to the predictions of KKMC, including test of the frames conventions. The plots of other elements $r_{i,j}$ are not as interesting, because they are very small (except for the $\tau$ mass effects), and therefore are not shown.
@@
The transverse spin correlations for $e^{-}e^{+}$ and $u\bar{u}$ change signs at energies below and above the $Z$-peak, and then for higher energies stabilize at rather small/large positive value, for $\theta=\pi/3$ and $\theta=2\pi/3$, respectively. This is not the case for $d\bar{d}$ initial state, where correlations change sign only below Z-boson peak, and then remain large and negative for $\theta=\pi/3$ or small and positive for $\theta=2\pi/3$, respectively. Above the $Z$-boson peak, they only very weakly depend on energy.
@@ 4 Reweighting procedure
The analytical formulas presented in Section 2 have been implemented into reweighting algorithms for KKMC and TauSpinner programs. Necessary tests were performed to confirm that event weights at fixed $\tau$-pair virtuality and incoming fermion flavor reproduce those formulas correctly.
@@ 4.1 The reweighting algorithm for KKMC: case of anomalous dipole moments
With respect to implementation prepared in Ref. [11], the function calculating the spin correlations and an overall event weight has been updated to include contributions due to $Z$-boson exchange and $Z\gamma$-interference contributions. The formulas have become substantially longer after introducing the anomalous couplings of the $\tau$ leptons to the $Z$-boson along with coupling to the virtual photon. Other technical aspects explained in Ref. [11] remain valid.
@@ 4.2 The reweighting algorithm for TauSpinner
The basic formalism of TauSpinner is documented in Eqs. (7) to (12) in Section 2.2 of Ref. [31]. Here we assume that reader is familiar with this reference, and on the details of how the kinematics of hard process can be deciphered from the information available in the event record.
@@
In this update, starting with the $q\bar{q}\to\tau^{-}\tau^{+}$ parton level processes, the functions calculating the spin correlations and an overall event weight have been updated in the TauSpinner with analytical formulas of Section 2. Before this update, the transverse spin correlations in process $q\bar{q}\to Z/\gamma^{*}\to\tau^{-}\tau^{+}$ were available via the the electroweak library SANC [32, 33] and its predefined tables, which were neither very flexible nor transparent.
Next, let us turn our attention to the new parton level process: $\gamma\gamma\to\tau^{-}\tau^{+}$. The formulas for spin amplitudes, cross section and spin-correlation matrix were presented in Section 2. Eqs. (29) contain all necessary expressions. It is important to note that each parton process contributes incoherently to the final state. Thus, the introduction of nearly on-shell photon as an incoming parton, and a corresponding hard process, is possible by straightforward extension of the sums in Eq. (33) only.
@@
In terms of technical details of the TauSpinner, we again assume that the reader is familiar with the ones described in Ref. [31]. Minor extensions to include additional parton level $\gamma\gamma$ process has been introduced in this update444 Practical comments can be found in the code of the latest version of TauSpinner, downloadable from the project web page http://tauolapp.web.cern.ch/tauolapp/..
Three new functions have been added to the TauSpinner code, and the sum in function double sigborn(...) was extended to the $\gamma\gamma$ process. In the function dipolgammarij(...), we calculate the $R_{i,j}^{\gamma\gamma}$ matrix. The functions:
T_gamm(MODE,SVAR,COSTHE,TA,TB) and T_gammNEW(MODE,SVAR,COSTHE,TA,TB) calculate cross-section normalized as in older parton level processes; anomalous dipole moments are absent in T_gamm() . The function T_gamm can be set to return zero with internal parameters of the algorithm. This feature is useful for reweighting samples which have no $\gamma\gamma$ contribution. The functions T_gamm and T_gammNEW are called by the algorithm with input parameters named MODE, SVAR, COSTHE, TA, and TB respectively. The MODE is not used at present. The Mandelstam $s$ variable SVAR, cosine of the hard process scattering angle COSTHE and helicities/chiralities TA, and TB of the outgoing $\tau$’s are set internally based on the event kinematics, and are not expected to be initialized by the user.
@@
  • iqed=iqedDip: to include SM magnetic dipole moment in $\gamma\gamma$ contribution in the denominator of the ratio defining weight, see Eqs. (19), (20) of Ref. [11]. The choice depends on whether the contribution is taken into account in the sample to be re-weighted. Note that this contribution is always present in the weight in the numerator, unless photon structure function is absent (set to zero).
@@
The implementation can be adapted for reweighting events of ultra-pheripheral heavy ions collisions, where it can be used for the analyses of $\tau$ dipole moments at LHC, as done in the Refs. [5, 6].
@@ 5 Numerical results for semi-realistic observable
In Ref. [11], we discussed the observables sensitive to contribution of the dipole form-factors in the $e^{-}e^{+}\to\gamma^{*}\to\tau^{-}\tau^{+}$ interaction at Belle II energies. We found that transverse spin correlations of the $\tau$-pair production, combined with $\tau^{\pm}\to\pi^{\pm}\pi^{0}\nu_{\tau}$ decays, may be useful for that purpose. For these energies, the transverse spin correlations, calculated at the lowest order in the SM are simple, and weakly dependent on the center-of-mass energy. Transverse spin correlations in the direction perpendicular to (aligned in) the reaction plane are of the opposite sign. Therefore, fortunately the corresponding kinematic configurations are easy to separate, and the initial-state bremsstrahlung (ISR) emissions do not contribute in sizable amounts to the transverse momenta of the $e^{-}e^{+}\to\tau^{-}\tau^{+}n\gamma$ events. The transverse momenta of unobservable photons $p_{T}^{ISR}$ are much smaller than $m_{\tau}$ and, as a consequence, also smaller than the transverse momenta of the $\tau$-decay products.
@@ 5.1 Observables sensitive to dipole moments at $Z$-boson pole
At the $Z$-boson peak, the spin-correlation pattern is different from that at low energies relevant for the Belle II experiments. The largest components of the $R$ matrix are of longitudinal-transverse type, e.g. the terms $R_{13}$, $R_{31}$,$R_{23}$,$R_{32}$ are large. That is why the observable of Ref. [11] is not sensitive to the anomalous magnetic and electric dipole form-factors.
We have chosen therefore different and also rather simple observable, for the case when both $\tau$ leptons decay to a $\pi$ and a $\nu_{\tau}$. In the center of mass of $\tau^{-}\tau^{+}$ pair, we take the energy to correspond to the mass of $Z$-boson. That should, in realistic conditions, reduce the size of the initial-state bremsstrahlung, even though the width of the $Z$-boson is larger than the mass of $\tau$-lepton. This aspect will require large care and detailed understanding of detector conditions in the future experiments at the Future Circular Collider (FCC-ee) at the $Z$-boson peak.

### v2

@@ 2.1 The $f_{i}\bar{f}_{i}\to\tau^{-}\tau^{+}$, $i=e,u,d$ case
$R_{33}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{2}}{8N_{i}D_{Z}(s)}\big{\{}16\beta\gamma^{2}\text{cos}(\theta)a_{\tau}\,a_{i}\,v_{\tau}\,v_{i}+\left(\gamma^{2}-1\right)$
@@
$R_{14}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{3}}{2N_{i}D_{Z}(s)}\text{sin}(\theta)\big{\{}\text{Im}(Y(s))\bigl{[}\beta\,\gamma^{2}\,v_{\tau}\,\text{cos}(\theta)(a_{i}^{2}+v_{i}^{2})$
@@
$R_{24}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{3}}{2N_{i}D_{Z}(s)}\text{sin}(\theta)\big{\{}\text{Im}(X(s))\bigl{[}\left(\gamma^{2}-1\right)v_{\tau}\,\text{cos}(\theta)(a_{i}^{2}+v_{i}^{2})$
@@
$R_{34}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{2}}{4N_{i}D_{Z}(s)}\big{\{}-4a_{i}\,v_{i}\,\text{cos}(\theta)(\left(\gamma^{2}-1\right)a_{\tau}^{2}+\gamma^{2}v_{\tau}^{2})$
@@
$R_{44}^{(Z)}=\frac{g_{Z}^{4}m_{\tau}^{4}\gamma^{2}}{8N_{i}D_{Z}(s)}\big{\{}16\beta\gamma^{2}\text{cos}(\theta)\,a_{\tau}\,a_{i}\,v_{\tau}\,v_{i}+\left(\gamma^{2}-1\right)(3+\text{cos}(2\theta))$
@@
  • Finally, the longitudinal polarization ${\cal P}_{L}=r_{34/43}$ of the $\tau^{-}$ and $\tau^{+}$ in the SM arises due to the vector coupling $v_{\tau}$. This is well known and ${\cal P}_{L}$ was measured already at LEP I, as noted in Ref. [31]. The longitudinal polarization, as can be seen from Eqs. (25), depends on the weak anomalous magnetic moment ${\rm Re}(X)$ as well.
@@
We would like to acknowledge that in case of $e^{-}e^{+}\to\tau^{-}\tau^{+}$ process, the terms linear in the spins and dipole moments were already derived in Refs. [30, 22]. However, our results represent some important extensions: they are obtained for various initial state fermions $i=(e,\,u,\,d)$, include the complete spin-correlation matrix between the decaying $\tau^{-}$ and $\tau^{+}$ leptons, and also include $Z$-boson exchange, $\gamma$ exchange and $\gamma\,Z$ interference.
@@ 2.2 The $f_{i}f_{i}\to\tau^{-}\tau^{+}$, $i=\gamma$ case
$R^{{\gamma\gamma}}_{13}$ $=$ $R^{\gamma\gamma}_{31}=\frac{e^{4}}{2D^{2}}\,\gamma\,\beta^{2}\big{[}(\beta^{2}+A\,(\beta^{2}-2)-1)\cos(2\theta)+A\beta^{2}-\beta^{2}+1\big{]}\,{\sin(2\theta)},$
@@
$R^{{\gamma\gamma}}_{23}$ $=$ $-R^{\gamma\gamma}_{32}=\frac{e^{4}B}{2D^{2}}\,\gamma\,\beta\,(\beta^{2}\cos(2\theta)-3\beta^{2}+2)\,{\sin(2\theta)},$
@@
All elements of $R^{\gamma\gamma}_{i,j}$, except the transverse-longitudinal ones, satisfy the condition $R^{{\gamma\gamma}}_{i,j}(\theta)=R^{{\gamma\gamma}}_{i,j}(\pi-\theta)$ for $0\leq\theta\leq\pi$.
@@
$+2(s^{-}_{2}s^{+}_{3}-s^{-}_{3}s^{+}_{2})\,\gamma(\beta^{2}\cos(2\theta)-3\beta^{2}+2)\,\sin(2\theta)\Big{]}.$
@@ 3 Transverse spin correlations of elementary processes
In this section, we present numerical results for $R_{i,j}$ for the process $f_{i}\,\bar{f}_{i}\to\tau^{-}\tau^{+}$ obtained with analytical formulas of Sec. 2, as a function of the invariant mass of $\tau$-lepton pair.
@@
Figure 1: Upper plots: transverse spin-correlation elements of the $r_{i,j}$ matrix: $r_{11}$ (solid lines) and $r_{22}$ (dashed lines) for the electron-positron initial state. The angle $\theta$ is chosen $\pi/3$ for the left plot and $32\pi/$ 3 for the right plot. The effective $Z$ couplings to leptons of Table 2 are used. Dipole moments are not included. Lower plots: $R_{44}/R_{44}(SM)$ at $\theta=\pi/3$ (left), and $\theta=2\pi/3$ (right). Solid lines: ${\rm Re}(A(s))=0.1$, dashed lines: ${\rm Re}(X(s))=0.1$, other form-factors are set to zero.
@@
We also used distributions from Fig. 2 to test our formulas with respect to the predictions of KKMC, including test of the frames conventions. The plots of other elements $r_{i,j}$ are not as interesting, because they are very small (except for the $\tau$ mass effects), and therefore are not shown.
@@
The transverse spin correlations for $e^{-}e^{+}$ and $u\bar{u}$ change signs at energies below and above the $Z$-peak, and then for higher energies stabilize at small positive value, except for $e^{-}e^{+}$ at $\theta=2\pi/3$, where value is larger. This is not the case for $d\bar{d}$ initial state, where correlations change sign only once below Z-boson peak, and then remain negative and small (large) for $\theta=\pi/3$ ($\theta=\pi/3$). Above the $Z$-boson peak, they only very weakly depend on energy.
In addition, in Fig. 2 (lower plots) we show dipole form-factors impact on the $R_{44}$ which determines the cross section of $e^{-}e^{+}\to\tau^{-}\tau^{+}$ of unpolarized $\tau$ leptons. The ratios of $R_{44}$ calculated with the form-factors ${\rm Re}(A(s))$ and ${\rm Re}(X(s))$ included, to $R_{44}(SM)$ calculated without form-factors, are presented. As is seen, there is a clear effect of the real parts of anomalous magnetic and anomalous weak magnetic form-factors. The electric form-factors $B(s)$ and $Y(s)$ do not contribute to $R_{44}$, while the contribution from ${\rm Im}(A(s))$ and ${\rm Im}(X(s))$ is very small and is not shown.
@@ 4 Reweighting procedure
The analytical formulas presented in Sec. 2 have been implemented into reweighting algorithms for KKMC and TauSpinner programs. Necessary tests were performed to confirm that event weights at fixed $\tau$-pair virtuality and incoming fermion flavor reproduce those formulas correctly.
@@ 4.1 The reweighting algorithm for KKMC: case of anomalous dipole moments
With respect to implementation prepared in Ref. [11], the function calculating the spin correlations and an overall event weight has been updated to include contributions due to $Z$-boson exchange and $Z\gamma$-interference. The formulas have become substantially longer after introducing the anomalous couplings of the $\tau$ leptons to the $Z$-boson along with coupling to the virtual photon. Other technical aspects explained in Ref. [11] remain valid.
@@ 4.2 The reweighting algorithm for TauSpinner
The basic formalism of TauSpinner is documented in Eqs. (7)–(12) in Sec. 2.2 of Ref. [34]. Here we assume that reader is familiar with this reference, and on the details of how the kinematics of hard process can be deciphered from the information available in the event record.
@@
In this update, starting with the $q\bar{q}\to\tau^{-}\tau^{+}$ parton level processes, the functions calculating the spin correlations and an overall event weight have been updated in the TauSpinner with analytical formulas of Sec. 2. Before this update, the transverse spin correlations in process $q\bar{q}\to Z/\gamma\to\tau^{-}\tau^{+}$ were available via the the EW library SANC [35, 28] and its predefined tables, which were neither very flexible nor transparent.
Next, let us turn our attention to the new parton level process: $\gamma\gamma\to\tau^{-}\tau^{+}$. The formulas for spin amplitudes, cross section and spin-correlation matrix were presented in Sec. 2. Eqs. (33) contain all necessary expressions. It is important to note that each parton process contributes incoherently to the final state. Thus, the introduction of nearly on-shell photon as an incoming parton, and a corresponding hard process, is possible by straightforward extension of the sums in Eq. (37) only.
@@
In terms of technical details of the TauSpinner, we again assume that the reader is familiar with the ones described in Ref. [34]. Minor extensions to include additional parton level $\gamma\gamma$ process has been introduced in this update555 Practical comments can be found in the code of the latest version of TauSpinner, downloadable from the project web page http://tauolapp.web.cern.ch/tauolapp/..
Three new functions have been added to the TauSpinner code, and the sum in function double sigborn(...) was extended to the $\gamma\gamma$ process. In the function dipolgammarij(...), we calculate the $R_{i,j}^{\gamma\gamma}$ matrix. The functions
T_gamm(MODE,SVAR,COSTHE,TA,TB) and T_gammNEW(MODE,SVAR,COSTHE,TA,TB) calculate cross section normalized as in older parton level processes; anomalous dipole moments are absent in T_gamm() . The function T_gamm can be set to return zero with internal parameters of the algorithm. This feature is useful for reweighting samples which have no $\gamma\gamma$ contribution. The functions T_gamm and T_gammNEW are called by the algorithm with input parameters named MODE, SVAR, COSTHE, TA, and TB respectively. The MODE is not used at present. The Mandelstam $s$ variable SVAR, cosine of the hard process scattering angle COSTHE and helicities/chiralities TA, and TB of the outgoing $\tau$’s are set internally based on the event kinematics, and are not expected to be initialized by the user.
@@
  • iqed=iqedDip: to include SM anomalous magnetic dipole moment in $\gamma\gamma$ contribution in the denominator of the ratio defining weight, see Eqs. (19), (20) of Ref. [11]. The choice depends on whether the contribution is taken into account in the sample to be reweighted. Note that this contribution is always present in the weight in the numerator, unless photon structure function is absent (set to zero).
@@
The implementation can be adapted for reweighting events of ultra-peripheral heavy-ion collisions, where it can be used for the analysis of $\tau$ dipole moments at the LHC, as done in the Refs. [5, 6].
@@ 5 Numerical results for semi-realistic observable
In Ref. [11], we discussed the observables sensitive to contribution of the dipole form-factors in the $e^{-}e^{+}\to\gamma^{*}\to\tau^{-}\tau^{+}$ interaction at Belle II energies. We found that transverse spin correlations of the $\tau$-pair production, combined with $\tau^{\pm}\to\pi^{\pm}\pi^{0}\nu_{\tau}$ decays, may be useful for that purpose. For these energies, the transverse spin correlations, calculated at the lowest order in the SM are simple, and weakly dependent on the c.m. energy. Transverse spin correlations in the direction perpendicular to (aligned in) the reaction plane are of the opposite sign. Therefore, fortunately the corresponding kinematic configurations are easy to separate, and the initial-state bremsstrahlung (ISR) emissions do not contribute in sizable amounts to the transverse momenta of the $e^{-}e^{+}\to\tau^{-}\tau^{+}n\gamma$ events. The transverse momenta of unobservable photons $p_{T}^{ISR}$ are much smaller than $m_{\tau}$ and, as a consequence, also smaller than the transverse momenta of the $\tau$-decay products.
@@ 5.1 Observables sensitive to dipole moments at $Z$-boson pole
At the $Z$-boson peak, the spin-correlation pattern is different from that at low energies relevant for the Belle II experiments. The largest components of the $R$ matrix are of longitudinal-transverse type, e.g. the terms $R_{13}$, $R_{31}$,$R_{23}$ and $R_{32}$ are large. That is why the observable of Ref. [11] is not sensitive to the anomalous magnetic and electric dipole form-factors.
We have chosen therefore different and also rather simple observable, for the case when both $\tau$ leptons decay to a $\pi$ and a $\nu_{\tau}$. In the center of mass of $\tau^{-}\tau^{+}$ pair, we take the energy to correspond to the mass of $Z$-boson. That should, in realistic conditions, reduce the size of the initial-state bremsstrahlung, even though the width of the $Z$-boson is larger than the mass of the $\tau$-lepton. This aspect will require large care and detailed understanding of detector conditions in the future experiments at the Future Circular Collider (FCC-ee) at the $Z$-boson peak.